OpenStudy (anonymous):

Use the definition of the Laplace to find the transform of f(t)=4^t. For what value of s is the transform defined? Hint: Rewrite the function in terms of e and the natural logarithm

OpenStudy (anonymous):

Use the hint: \[\large 4^t=e^{\ln4^t}=e^{t\ln 4}\]